HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THEORETICAL STUDY OF THE ANOMALOUS BAND INTENSITY PROFILE OF THE ν_{3} + ν_{5} BAND OF SF_{6}

\underline{M. LOËTE} , \underline{V. BOUDON}, Laboratoire Interdisciplinaire Carnot de Bourgogne,

UMR 6303 CNRS-Université Bourgogne, 9 Avenue Alain Savary, BP 47 870, F-21078 DIJON Cedex FRANCE; \underline{L. MANCERON} , AILES Beamline, Synchrotron

SOLEIL, L’Orme des Merisiers, BP 48 F-91192, SAINT-AUBIN, FRANCE

As part of an extensive study of the high-resolution infrared spectroscopy of the SF_{6} molecule, Faye et al. have revealed the very unusual band intensity profile of the ν_{3} + ν_{5} combination band^{1}. This one is very asymmetric and consists of a huge P branch along with weak Q and R branches (see Figure below).

In order to explain this structure, we undertook a theoretical study about the development of the effective dipole moment operator μ̃ for the ν_{3} + ν_{5} band considered as isolated, using the contact transformation approach. Starting from the Hamiltonian operator, we have diagonalized the first order (H_{1}) terms that contain, in particular, Coriolis contributions. From this, we could deduce the expression of μ̃ via the first contact transformation. The experimental parameter values coincide well with calculations made using these expressions. The small differences can be explained by higher order terms (second contact transformation).

This study also allows to estimate the zero order parameter of the untransformed dipole moment operator for the ν_{4} + ν_{5} and ν_{5} + ν_{6} combination bands.

Experimental vs. simulated spectrum of the ν_{3} + ν_{5} band of SF_{6}.

1M. Faye, L. Manceron, P. Roy, V. Boudon and M. Loëte, J. Mol. Spectrosc. 348, 37–42 (2018).

p-number: p032  Submitted on Mon, 05 Jun 2023 10:03:38 +0200